Circular No. 5934

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

A0535+26 M. H. Finger, Computer Sciences Corporation; and R. B. Wilson

and B. A. Harmon, NASA/Marshall Space Flight Center, report for the BATSE/Compton GRO Observatory team: “The hard x-ray flux currently being observed from the x-ray binary pulsar A0535+26 (IAUC 5931, 5933) has significant aperiodic variability in addition to coherent pulsations. Power spectra contain the pulsed component and an underlying continuum with a significant bump between approximately 20 and 80 mHz, in both the 20–50 and 50–100-keV energy bands. The continuum in the 20–50-keV power spectra on Feb. 9 and 12 has been fitted with a sum of power-law and gaussian components. On both days the power-law component has an index of ∼ 1.0 and a fractional RMS (10–500 mHz) ∼ 20 percent. On Feb. 9 the bump was centered at 50 ± 1 mHz, had an FWHM of 28 ± 3 mHz and a fractional RMS of 14.1 ± 0.5 percent. On Feb. 12 the bump was centered at 65 ± 1 mHz, had a FWHM of 37 ± 3 mHz and a fractional RMS of 11.6 ± 0.4 percent. The 20–40-keV flux increased from 3.1 ± 0.2 Crab on Feb. 9 to 5.6 ± 0.2 Crab on Feb. 12. The shift in center frequency is therefore consistent with the luminosity dependence predicted by the beat-frequency QPO model (Alpar and Shaham 1985, Nature 316, 239). Both the total and pulsed flux have continued to increase. On Feb. 14 the 20–40-keV total flux was 6.2 ± 0.1 Crab, and the phase-averaged pulsed flux was 1.6 ± 0.1 Crab (total).”

NOVA CASSIOPEIAE 1993 Recent visual magnitude estimates indicate that the nova is now fading

rapidly: Feb. 1.13 UT, 8.0 (J. Pleszka, Cracow, Poland); 2.12, 8.3 (C. E. Spratt, Victoria, BC); 3.09, 8.4 (J. D. Shanklin, Cambridge, England); 4.72, 8.5 (A. Lauvstad, Levanger, Norway); 5.86, 8.3 (P. Rapavý, Rimavská Sobota, Slovakia); 6.76, 8.2 (M. Verdenet, Bourbon-Lancy, France); 8.02, 8.2 (G. W. Kronk, Troy, IL); 9.11, 8.1 (R. Royer, Lakewood, CA); 10.23, 8.3 (P. Schmeer, Bischmisheim, Germany); 12.03, 8.5 (R. Hays, Worth, IL); 12.80, 8.4 (B. H. Granslo, Fjellhamar, Norway); 13.10, 8.6 (Royer); 13.73, 9.0 (Granslo); 14.76, 9.0 (Schmeer); 15.06, 9.2 (W. G. Dillon, Missouri City, TX); 15.78, 9.6 (Granslo); 16.04, 9.9 (Hays); 16.78, 10.1 (Schmeer); 17.06, 10.1 (Dillon); 17.77, 10.5 (T. Vanmunster, Landen, Belgium).

1994 February 17 Brian G. Marsden